OpenStudy (anonymous):

Pipes A and B can fill a storage tank in 8 hours and 12 hours, respectively. With the tank empty, pipe A was turned on at noon, and then pipe B was turned on at 1:30 P.M. At what time was the tank full?

OpenStudy (anonymous):

no one likes these questions i see

OpenStudy (amistre64):

this is an easier one tho :)

OpenStudy (anonymous):

ok an hour and a half for pipe A that fill the tank in 8 hours, means its rate is 1/8 per hour and \[1.5\times \frac{1}{8}=\frac{3}{16}\] meaning the tank is \[\frac{3}{16}\] full at one thirty. the two pipes together work at a rate of \[\frac{1}{8}+\frac{1}{12}=\frac{5}{24}\] of the tank per hour. now solve \[\frac{3}{16}+\frac{5}{24}x=1\] for x

OpenStudy (anonymous):

did other annoying refund one too, but i think they have left already

OpenStudy (amistre64):

yeah, its a thankless job eh ;)

OpenStudy (anonymous):

did i do this right because i get 3.9 usually problems cooked to nice integer solutions